# MedFact: Benchmarking the Fact-Checking Capabilities of Large Language Models on Chinese Medical Texts

### v1

@@ Introduction
Table 2: Performance of DeepSeek-R1 and XiaoYi across different strategies on MedFact. For each model, the best score is in bold and the second-best is underlined.
Main Results
Overall Performance Table 1 summarizes the performance of all evaluated models. We report the F1 score for VC and Precision, Recall, and F1 for EL. Among the models evaluated, proprietary models consistently outperform their open-source counterparts across both tasks and all prompting strategies. Doubao-Seed-1.6-thinking emerges as the top-performing general-purpose model, while XiaoYi leads among medically specialized models. A significant performance gap remains between the best-performing models and human experts. The top model F1 score on the EL task, 0.6858, achieved by XiaoYi with CoT, is still below the human baseline of 0.7012. This disparity underscores both the difficulty of MedFact and the substantial challenge for current LLMs. The full results are provided in the Appendix.
A comparison between the tasks reveals a consistent performance gap, with all models performing worse on EL than on VC. For instance, the F1 scores for DeepSeek-V3 and Qwen2.5-Max decrease from 0.6947 and 0.7352 on VC to 0.5380 and 0.6116 on EL, respectively. This suggests that while models can often determine if a text contains an error, they struggle to pinpoint where it is. This difficulty often arises from a “correct-for-the-wrong-reason” phenomenon: a model may correctly classify a text as incorrect (a successful VC judgment) but do so by misidentifying a factually sound statement as the source of the error. Such behavior, rooted in a lack of deep medical knowledge, demonstrates that the EL task is a more stringent measure of a model’s fine-grained understanding and reveals fundamental deficiencies in its underlying medical knowledge.
Impact of Advanced Reasoning Strategies The results of DeepSeek-R1 and XiaoYi with advanced reasoning strategies against a zero-shot baseline are presented in Table 10. The retrieval corpus for RAG consists of the 6,405 expert-annotated source texts from our data curation process, excluding any rewrites. Results show that RAG using either one or three retrieved references yields substantial performance gains on both the VC and EL tasks. This finding shows that providing models with direct access to external, domain-specific knowledge is a highly effective method for medical fact-checking. MedPrompt delivers modest improvements, demonstrating that performance can also be enhanced solely through prompting techniques such as few-shot learning and self-consistency.
In contrast, the multi-agent frameworks (MAD and MDAgents) exhibit a consistent trade-off: they improve Recall at the expense of Precision, resulting in a stable F1 score for MAD and only a marginal improvement for MDAgents. We attribute this pattern to a phenomenon we term “over-criticism”. The collaborative process, which encourages the generation of multiple perspectives and error hypotheses, appears to increase the models’ propensity to flag errors where none exist, thereby elevating the false positive rate and lowering Precision. The slight performance edge of MDAgents may be attributed to its use of medical role-play, which could temper this critical tendency. This finding reveals a critical challenge: while these methods are designed to improve reasoning, they can inadvertently encourage an overly skeptical posture. This undermines reliability in medical domains, where correctly identifying accurate information is as crucial as flagging falsehoods.
@@
Analysis
Inference-time Scaling
Motivated by the superior performance of inference-time reasoning models such as DeepSeek-R1 and Doubao-Seed-1.6-thinking, over their models from the same series (DeepSeek-V3 and Doubao-Seed-1.6), we investigate the potential of inference-time scaling strategies. Specifically, we evaluate the impact of budget forcing (Muennighoff et al. 2025). This technique modulates a model’s reasoning process at inference time by either prematurely terminating generation or prolonging it via “Wait” tokens to adjust the computational budget. We focus on two models adapted for this paradigm: s1.1-32B and m1-32B-1K. s1.1-32B is a general reasoning model fine-tuned from Qwen2.5-32B-Instruct on only 1,000 high-quality reasoning examples (Muennighoff et al. 2025). m1-32B-1K is a specialized medical reasoning model fine-tuned from Qwen2.5-32B-Instruct, using 1,000 medical question-answering pairs (Huang et al. 2025).
Table 11 compares the performance of the baseline model (Qwen2.5-32B-Instruct) against s1.1-32B and m1-32B-1K under a CoT setting, evaluated both with and without budget forcing. The results indicate that s1.1-32B, fine-tuned on general-purpose data, suffers a substantial performance degradation on the EL task, which we attribute to its lack of domain-specific medical knowledge due to fine-tuning. In contrast, m1-32B-1K, with its medical fine-tuning, outperforms the base model on both tasks. However, applying budget forcing yields no performance gain and instead exacerbates the “over-criticism” phenomenon. The prolonged reasoning process causes models to hallucinate errors, a behavior analogous to the “over-thinking” observed in reasoning models (Ghosal et al. 2025; Sui et al. 2025). This suggests that simply extending a model’s reasoning time without providing new information can lead to counterproductive deliberation rather than improved accuracy. Ultimately, this finding reinforces the conclusion that for knowledge-intensive tasks like medical fact-checking, enhancing a model’s underlying medical knowledge is more critical than simply increasing its computational budget at inference.
@@
Error Analysis
To better understand the performance gap between LLMs and human experts, and to diagnose why the EL task is particularly challenging, our human experts manually annotated and categorized the errors made by the top-performing models. As illustrated in Figure 4, the lack of medical knowledge constitutes the predominant category of model mistakes (76.09% of all errors for DeepSeek-R1), where models demonstrate insufficient understanding of medical principles, conflating distinct conditions or treatments, and relying on outdated information, which indicates that models often struggle with the nuanced, specialized knowledge required for accurate medical fact-checking.
Fine-grained Performance
We conduct a fine-grained analysis of model performance across different writing styles (Figure 9). For the VC task, models are more effective at identifying errors in fabricated medical misinformation than in texts from encyclopedias. This indicates that models can readily detect exaggerated falsehoods but struggle with the subtle, nuanced inaccuracies often present in professionally curated content. Notably, this performance gap diminishes for the more demanding EL task. This shift suggests that performance on localization is less influenced by stylistic cues and more fundamentally dependent on the model’s underlying medical knowledge. Consequently, the EL task serves as a more direct and robust measure of a model’s actual domain expertise.
@@
Data Contamination
To investigate potential data contamination within the pre-training corpora of the evaluated LLMs, we devised a text completion task. For each instance in MedFact, we prompted the models with the first half of the text and tasked them with generating the continuation. We then measured the textual overlap between the generated output and the ground-truth second half using ROUGE-F1 scores, where a high score would suggest memorization. As shown in Table 12, all evaluated models, even those that achieve state-of-the-art performance on our tasks, exhibit negligible ROUGE scores. This result strongly indicates that MedFact serves as a valid and previously unseen benchmark for these models.
Conclusion
@@
We introduce MedFact, a new benchmark for evaluating the fact-checking abilities of LLMs in Chinese medical texts. Constructed via a meticulous hybrid AI-human pipeline, our dataset is diverse, realistic, and uncontaminated, allowing it to probe for nuanced medical understanding. Our comprehensive evaluation of 20 leading LLMs reveals a critical performance deficit compared to human experts, particularly on the fine-grained task of error localization. Furthermore, LLMs exhibit an “over-criticism” failure mode that significantly degrades their precision. The challenges highlighted underscore the substantial work required to develop LLMs that are not just knowledgeable but also factually reliable enough for safe deployment in medical applications.
Appendix A Ethics Statement
Human Annotation and Compensation All expert annotators who contributed to MedFact received fair compensation for their time and expertise at rates that meet or exceed the minimum wage in their respective regions.
Privacy and Consent We employed a rigorous anonymization process to remove all Personally Identifiable Information (PII) and ensure privacy protection. Our use of this data is strictly for non-commercial research purposes and complies with the terms of service of source platforms under fair use principles.
Risk of Misinformation The factually incorrect instances in MedFact could be exploited to train models that generate harmful content or to disseminate misinformation directly. We will distribute the dataset under a restrictive license that explicitly prohibits use in clinical or patient-facing applications. The dataset will be accompanied by comprehensive documentation that clearly warns users about the presence of factual errors and specifies that its intended purpose is exclusively for research and benchmarking applications.
Bias and Representativeness MedFact focuses exclusively on Chinese medical texts. Consequently, the data, error distributions, and model findings may not generalize to other languages, cultures, or healthcare systems. We caution researchers against drawing overly broad conclusions from our findings and encourage the development of analogous benchmarks for other languages and cultural contexts.
@@ Appendix B Dataset Construction & Statistics
Multi-criteria Filtering
Across the three filtering rounds, the acceptance rates (ratio of passed instances to candidate pool size) were 67.69%, 37.00%, and 23.62%, respectively. The consistency between the model suite’s judgment and medical professionals increased progressively: 69.72%, 84.37%, and 96.40%, respectively, while the reference database size grew from 412 to 624 instances. Figure 10 presents one accepted example, and Figure 11 presents three rejected examples from the reference database during the multi-criteria filtering process, illustrating the filtering criteria: overly simplistic, esoteric, and malformed content.
Error Taxonomy Definitions
@@
Non-Medical Errors Non-medical errors include factual inaccuracies that are not domain-specific to medicine.
@@
Writing Style Taxonomy Definitions
@@ II. Non-Medical Errors
Data Examples
@@ Appendix C Implementation Details
Prompt Templates
Figures 12 to 13 present the prompts used for zero-shot and CoT prompting. All prompts were originally formulated in Chinese; we provide their corresponding English translations for clarity and reference.
Model Versions
We accessed proprietary models through API calls and deployed open-source models locally. Table 5 lists the model versions and full names used in our experiments.
LLM-as-a-Judge
For the evaluation of error localization tasks, we assessed the consistency agreement between different models and human experts. DeepSeek-R1, GPT-4o, o1, Claude 3.7 Sonnet, and Gemini 2.5 Pro achieved Cohen’s Kappa values of 0.800, 0.870, 0.870, 0.845, and 0.865, respectively, on a random 10% sample of the veracity classification task results from Doubao-Seed-1.6-thinking, XiaoYi, and Qwen-Max. Considering API costs and the evaluation performance, we selected GPT-4o for the LLM-as-a-judge evaluation.
Appendix D Complete Experimental Results
Full Result Tables
@@ Appendix D Complete Experimental Results
Fine-grained Performance
Figure 9 details the fine-grained model performance across various error types, medical specialties, and writing styles.

### v2

@@ 4.1 Experimental Setup
Table 2: Performance of DeepSeek-R1 and XiaoYi across different strategies on MedFact.
4.2 Main Results
Overall Performance Table 1 summarizes the evaluated model performance. Proprietary models consistently outperform open-source counterparts across both tasks and all prompting strategies. Doubao-Seed-1.6-thinking emerges as the top general-purpose model, while XiaoYi leads among medically specialized models. A significant performance gap remains between best-performing models and human experts. The top model F1 score on EL, 0.6858, achieved by XiaoYi with CoT, falls below the human baseline of 0.7012.
Task comparison reveals that all models perform worse on EL than on VC. For instance, F1 scores for DeepSeek-V3 and Qwen2.5-Max decrease from 0.6947 and 0.7352 on VC to 0.5380 and 0.6116 on EL, respectively. This suggests that while models can often determine if text contains errors, they struggle to pinpoint the location. This difficulty often arises from a “correct-for-the-wrong-reason” phenomenon: models may correctly classify text as incorrect but misidentify factually sound statements as error sources. Such behavior, rooted in a lack of medical knowledge, demonstrates that EL is a more stringent measure of fine-grained understanding and reveals fundamental deficiencies in underlying medical knowledge.
Advanced Reasoning Strategies Table 2 presents DeepSeek-R1 and XiaoYi results with advanced reasoning strategies against zero-shot baselines. The RAG corpus consists of 6,405 expert-annotated source texts from our data curation process. Results show RAG yields substantial performance gains on both VC and EL tasks. This demonstrates that providing models with direct access to external, domain-specific knowledge is highly effective for medical fact-checking. In contrast, multi-agent frameworks exhibit a trade-off: they improve Recall at the expense of Precision, resulting in stable F1 for MAD and marginal improvement for MDAgents. We attribute this to “over-criticism”: the collaborative process encourages generating multiple perspectives and error hypotheses, which increases models’ propensity to flag errors where none exist, thereby elevating false positive rates and lowering Precision. This reveals a challenge that, while these methods are designed to improve reasoning, they can encourage overly skeptical postures in medical fact-checking.
@@ 4.2 Main Results
5 Analysis
5.1 Inference-time Scaling
We evaluate budget forcing (muennighoff2025s1simpletesttimescaling), which modulates model reasoning at inference time by either prematurely terminating generation or prolonging it via “Wait” tokens to adjust computational budget. We focus on two models adapted for this paradigm: s1.1-32B muennighoff2025s1simpletesttimescaling and m1-32B-1K. s1.1-32B is a general reasoning model fine-tuned from Qwen2.5-32B-Instruct on only 1,000 high-quality reasoning examples huang2025m1unleashpotentialtesttime. m1-32B-1K is a specialized medical reasoning model fine-tuned from Qwen2.5-32B-Instruct, using 1,000 medical question-answering pairs. Table 3 compares the performance of the baseline model against s1.1-32B and m1-32B-1K under a CoT setting. The results indicate that s1.1-32B, fine-tuned on general-purpose data, suffers a substantial performance degradation on the EL task, which we attribute to its lack of domain-specific medical knowledge. In contrast, m1-32B-1K, with its medical fine-tuning, outperforms the base model. Applying budget forcing yields no performance gain and instead exacerbates the “over-criticism” phenomenon. The prolonged reasoning process causes models to hallucinate errors, a behavior analogous to the “over-thinking” observed in reasoning models ghosal2025doesthinkinghelpunderstanding. This suggests that simply extending a model’s reasoning time without providing new information can lead to counterproductive deliberation, and reinforces the conclusion that for medical fact-checking tasks, enhancing a model’s underlying medical knowledge is more critical.
@@ 5.1 Inference-time Scaling
5.2 Error Analysis
Human experts manually annotated and categorized top-performing model errors. As illustrated in Figure 4, lacking medical knowledge constitutes the predominant error category (76.09% of all errors for DeepSeek-R1), where models demonstrate insufficient understanding of medical principles, conflate distinct conditions or treatments, and rely on outdated information, indicating that models often struggle with nuanced, specialized knowledge required for accurate medical fact-checking.
5.3 Fine-grained Performance
We conduct fine-grained analysis of model performance across different writing styles (Figure 5). For VC, models are more effective at identifying errors in fabricated medical misinformation. This indicates that models can readily detect exaggerated falsehoods but struggle with subtle, nuanced inaccuracies often present in professionally curated content, and suggests that localization performance is less influenced by stylistic cues and more fundamentally dependent on underlying medical knowledge. This observation shows that EL serves as a more robust measure of medical expertise.
5.4 Data Contamination
To investigate potential data contamination within the pre-training corpora of the evaluated LLMs, we devised a text completion task. For each instance in MedFact, we prompted the models with the first half of the text and tasked them with generating the continuation. We then measured the textual overlap between the generated output and the ground-truth second half using ROUGE-F1 scores, where a high score would suggest memorization. As shown in Table 4, all evaluated models, even those that achieve state-of-the-art performance on our tasks, exhibit negligible ROUGE scores. This result indicates that MedFact serves as a valid and previously unseen benchmark for these models.
@@ 5.4 Data Contamination
6 Conclusion
@@
We introduce MedFact, a new benchmark for evaluating the fact-checking abilities of LLMs in Chinese medical texts. Constructed via a hybrid AI-human pipeline, our dataset is diverse, realistic, and uncontaminated, allowing it to probe for nuanced medical understanding. Our comprehensive evaluation of 20 leading LLMs reveals a critical performance deficit compared to human experts, particularly on the fine-grained task of error localization. Furthermore, LLMs exhibit an “over-criticism” failure mode that significantly degrades their precision. The challenges highlighted underscore the substantial work required to develop LLMs that are not just knowledgeable but also factually reliable enough for safe deployment in medical applications.
Limitations
MedFact targets the Chinese language within China’s healthcare system, and performance may not generalize to other languages, cultures, or healthcare systems. This focus ensures high linguistic quality and cultural consistency, though future work could develop parallel benchmarks for cross-lingual comparison. For the EL task, we use GPT-4o as an automated judge, which enables scalable evaluation but has inherent limitations, including potential biases and inconsistencies. Future work could explore hybrid evaluation frameworks combining LLM judge scalability with expert validation for ambiguous or high-stakes cases. Additionally, MedFact captures medical knowledge at creation time, and as medical science evolves, some information may become outdated. We prioritize foundational medical knowledge to minimize this issue, but acknowledge that dynamic benchmark updates would be valuable in the long term.
Ethical Considerations
All expert annotators received fair compensation at or above minimum wage in their regions. We applied anonymization to remove all Personally Identifiable Information (PII), and data use is strictly for non-commercial research and complies with source platform terms under fair use principles. Factually incorrect instances could be exploited to train harmful content generators or disseminate misinformation, so we distribute the dataset under a restrictive license prohibiting clinical or patient-facing applications. The documentation clearly warns about factual errors and specifies that the dataset is intended exclusively for research and benchmarking purposes. Additionally, MedFact focuses exclusively on Chinese medical texts, and data, error distributions, and findings may not generalize to other languages, cultures, or healthcare systems. We caution against drawing overly broad conclusions and encourage the development of analogous benchmarks for other languages and contexts.
Appendix A Detailed Related Works
A.1 LLMs in Medical Domains
The capabilities of Large Language Models (LLMs) across many disciplines have prompted the exploration of their applications in the medical domain huang2024comprehensivesurveyevaluatinglarge. Medprompt nori2023generalistfoundationmodelsoutcompete enhances LLMs’ performance on medical benchmarks by combining self-generated Chain-of-Thought (CoT) NEURIPS2022_9d560961, few-shot learning with dynamic example selection liu-etal-2022-makes, and self-consistency wangself. Other developments include HuaTuo wang2023huatuotuningllamamodel, a LLaMA-based model fine-tuned on Chinese medical instruction data; HuatuoGPT zhang-etal-2023-huatuogpt, a model specialized for medical consultation, fine-tuned on responses distilled from ChatGPT and real-world data; and PediatricsGPT NEURIPS2024_fa5b423e, a model developed via a systematic training pipeline for pediatric and general medical applications. These advances highlight a growing interest in adapting LLMs to medical reasoning and the specialized requirements of real-world practice.
A.2 Benchmarking Medical Knowledge in LLMs
Numerous benchmarks have been developed to evaluate the medical knowledge of LLMs nguyen-etal-2023-medredqa; NEURIPS2024_e15c4aff; abacha2025medecbenchmarkmedicalerror. MedQA app11146421 is a multiple-choice QA dataset derived from the United States Medical Licensing Examinations to assess the clinical knowledge of LLMs. Huatuo-26M wang-etal-2025-huatuo is a Chinese dataset comprising data from online consultation websites, encyclopedias, and information extracted from knowledge bases. More recently, HealthBench arora2025healthbenchevaluatinglargelanguage is a benchmark of realistic health conversations evaluated with custom, physician-created rubrics. MedXpertQA zuo2025medxpertqabenchmarkingexpertlevelmedical evaluates expert-level medical knowledge using both textual and multimodal tasks.
While some benchmarks for medical fact-checking exist, they often focus on narrow contexts. VeriFact Chung2025 verifies synthetic text against structured Electronic Health Records using a retrieval-augmented generation (RAG) and LLM-as-a-Judge paradigm. MEDEC abacha2025medecbenchmarkmedicalerror provides a benchmark for error detection and correction, but is confined to clinical notes. Relying on synthetic data or a single text genre, these benchmarks fail to capture the diverse contexts of medical information encountered in the real world, thereby limiting their utility for assessing how well models generalize. Consequently, a gap remains for a benchmark that can robustly evaluate fact-checking across a wide array of realistic medical texts.
A.3 Benchmarking Factuality in LLMs
Recent research has produced several benchmarks to evaluate factuality in LLMs. For instance, HaluEval li-etal-2023-halueval focuses on hallucination detection through a large-scale collection of automatically generated and human-annotated examples. SimpleQA wei2024measuringshortformfactualitylarge measures short-form factuality using single-answer questions adversarially collected to challenge state-of-the-art models like GPT-4. More broadly, OpenFactCheck iqbal-etal-2024-openfactcheck proposes a unified framework for benchmarking fact-checking across diverse sources, including free-form text and LLM-generated content. While these benchmarks provide valuable tools for assessing general-domain factuality, they lack the domain-specific nuances required for high-stakes fields like medicine, where misinformation can have severe public health consequences 10.1001/jamapediatrics.2023.5282. Addressing this challenge requires a solution that bridges two critical gaps: the lack of domain-specific context in general factuality benchmarks and the lack of real-world diversity in existing medical fact-checking benchmarks. To this end, we introduce MedFact, a new benchmark designed to evaluate the fact-checking capabilities of LLMs on a comprehensive and diverse collection of realistic medical texts.
@@ Appendix B Dataset Construction & Statistics
B.1 Multi-criteria Filtering
Across the three filtering rounds, the acceptance rates (ratio of passed instances to candidate pool size) were 67.69%, 37.00%, and 23.62%, respectively. The consistency between the model suite’s judgment and medical professionals increased progressively: 69.72%, 84.37%, and 96.40%, respectively, as the reference database size grew from 412 to 624 instances. Figure 10 presents one accepted example, and Figure 11 presents three rejected examples from the reference database during the multi-criteria filtering process, illustrating the filtering criteria: overly simplistic, esoteric, and malformed content.
B.2 Error Taxonomy Definitions
@@
Non-Medical Errors Non-medical errors encompass factual inaccuracies that are not specific to the medical domain.
@@
B.3 Writing Style Taxonomy Definitions
@@ B.4 Data Statistics
B.5 Data Contamination
To investigate potential data contamination within the pre-training corpora of the evaluated LLMs, we devised a text completion task. For each instance in MedFact, we prompted the models with the first half of the text and tasked them with generating the continuation. We then measured the textual overlap between the generated output and the ground-truth second half using ROUGE-F1 scores, where a high score would suggest that the model had memorized the content. As shown in Table 14, all evaluated models exhibit negligible ROUGE scores, which strongly indicates that MedFact serves as a previously unseen benchmark.
B.6 Data Examples
@@ Appendix C Implementation Details
C.1 Prompt Templates
Figures 12 to 13 present the prompts used for zero-shot and CoT prompting. All prompts were originally formulated in Chinese; we provide their corresponding English translations for reference.
C.2 Model Versions
We accessed proprietary models through API calls and deployed open-source models locally. Table 5 lists the model versions and full names.
C.3 LLM-as-a-Judge
For the evaluation of EL tasks, we assessed the agreement between different models and human experts. DeepSeek-R1, GPT-4o, o1, Claude 3.7 Sonnet, and Gemini 2.5 Pro achieved Cohen’s Kappa values of 0.800, 0.870, 0.870, 0.845, and 0.865, respectively, on a random 10% sample of the veracity classification task results from Doubao-Seed-1.6-thinking, XiaoYi, and Qwen-Max. Considering API costs and evaluation performance, we selected GPT-4o for the LLM-as-a-judge evaluation.
Appendix D Experimental Results
D.1 Full Result Tables
@@
Model Precision Recall F1 Score
DeepSeek-R1 0.5035 0.7580 0.6051
   + RAG (top-3 from same sources) 0.6112 0.7713 0.6820
   + RAG (top-3 from authoritative sources) 0.5557 0.7069 0.6223
XiaoYi 0.6512 0.7023 0.6758
   + RAG (top-3 from same sources) 0.6985 0.7117 0.7051
   + RAG (top-3 from authoritative sources) 0.6861 0.6796 0.6828
Table 12: Performance of different models with different sources for RAG on MedFact using zero-shot prompting. The best results are bolded and the second-best are underlined.
Model Language Precision Recall F1 Score
Gemini 2.5 Pro Chinese 0.4828 0.8752 0.6223
English 0.5286 0.9317 0.6745
Doubao-Seed-1.6-thinking Chinese 0.6307 0.7344 0.6786
English 0.5952 0.8483 0.7000
Table 13: Performance comparison between Chinese and English versions of MedFact using CoT prompting. The best results are bolded and the second-best are underlined.
@@
D.2 Impact of RAG Reference Sources
To evaluate how corpus selection affects medical fact-checking performance, we performed experiments comparing reference texts from authoritative sources (including peer-reviewed journals, clinical guidelines, and medical literature) with generic sources. Using the same retrieval pipeline, we assessed the impact of reference text relevance on fact-checking outcomes. As shown in Table 12, while precision improves slightly compared to zero-shot baselines when using authoritative sources, performance remains substantially lower than with highly relevant references. Crucially, recall decreases below the zero-shot baseline when using generic authoritative texts. We attribute this to relevance mismatch: irrelevant references severely degrade both precision and recall by biasing models toward accepting unverified content as correct, which shows a critical risk for industrial RAG deployments where reference quality varies.
D.3 Cross-lingual Evaluation on English Translation
To assess the impact of language on model performance, we evaluated an English-translated version of MedFact. We selected two models: Doubao-Seed-1.6-thinking and Gemini 2.5 Pro. Table 13 compares performance between Chinese and English versions using CoT prompting. Both models show F1 improvements on English: Doubao-Seed-1.6-thinking improves from 0.6786 to 0.7000, while Gemini gains more substantially from 0.6223 to 0.6745. However, both models consistently show decreased Precision and increased Recall in English. This indicates the over-criticism phenomenon persists and intensifies across languages. Models flag potential errors more aggressively in English, producing higher false positive rates despite improved F1 scores. These findings show language choice affects both overall performance and the precision-recall trade-off, with important implications for multilingual deployment in healthcare organizations.
D.4 Fine-grained Performance
We conduct fine-grained analysis of model performance across various error types, medical specialties, and writing styles (Figure 9). Across different writing styles, for VC, models are more effective at identifying errors in fabricated medical misinformation than in encyclopedia texts. This suggests that models can readily detect exaggerated falsehoods but struggle with subtle, nuanced inaccuracies that are often present in professionally curated content. This suggests that localization performance is less influenced by stylistic cues and depends more fundamentally on underlying medical knowledge. Consequently, EL serves as a more robust measure of actual domain expertise.